# ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze

### v2

@@ 4.2.1 Reuse the Information from Backward
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform the search in a reverse order. That is to say, we first perform tree search on the state of the last time-step in the trajectory, and then search the second to last, and so on. After we search on state $s_{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes. (the state of the $t+1$-th time-step in the trajectory), the root value $m_{t+1}$ is obtained. When performing search on $s_{t}$, we set the value of $s_{t+1}$ to the fixed value $m_{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $s_{t}$ according to the following equation:
$a_{root}=\arg\max_{a}I_{t}(a)$ (9)
$I_{t}(a)=\begin{cases}UCB_{score}(s_{t},a),&a\neq a_{t}\\ r_{t}+\gamma m_{t+1},&a=a_{t}\end{cases}$ (10)
where $a$ is an action corresponding to a child node, $a_{t}$ is the action corresponding to $s_{t+1}$, and $r_{t}$ is the reward predicted by the dynamic model. If an action other than $a_{t}$ is selected, the simulation just traverses following the original setting in MCTS (Schrittwieser et al., 2019). If action $a_{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $s_{t+1}$ is saved, the modified search process is faster than the original version, achieving Speedy Reanalyze.
@@
4.2.2 Theoretical Analysis
AlphaZero employs a variant of the UCB algorithm during MCTS search. Subsequently, MuZero proposes further improvements to this variant formula. Practical evidence has demonstrated the effectiveness of both these node selection methods; however, they have not been theoretically analyzed in their original paper. In earlier studies, UCT analyzed the convergence properties of the UCB1 algorithm in a tree structure, albeit without introducing a prior policy. PUCB (Rosin, 2011) incorporates relevant analysis of a prior policy, but it uses the prior in a penalty function-like approach which is different from in the MCTS-based algorithm family. In recent work, Gumbel MuZero (Danihelka et al., 2022) approximated the tree search model as a stochastic bandit, and on this basis, proved the monotonic improvement of their designs. In this section, we will draw upon such a simplification and demonstrate within the context of a stochastic bandit model why the node selection formula and policy updating methods used in the MCTS-based algorithm family are effective, and how information reuse can influence the policy target.
Definition 4.1.
After $n$ simulations with a stochastic bandit $\nu$ with $K$ arms, let $T_{i}$ denote the number of visits to arm $i$, then the policy generated based on the visit distribution is a random vector $P=[\frac{T_{1}}{n},...,\frac{T_{K}}{n}]$ The value $V_{\nu}(P)$ of policy $P$ on $\nu$ is defined as:
$V_{\nu}(P)=\sum_{i}\frac{T_{i}\mu_{i}(\nu)}{n}$ (11)
In cases where it does not cause any ambiguity, we will omit the mark $\nu$ in the following text. The next theorem explains the rationale behind using distribution as a target policy:
Theorem 4.2.
If an algorithm $\pi$ exhibits sub-linear growth in regret $R_{n}$ with respect to a bandit $\nu$, then the expectation $E[V(P)]$ converges to $\mu^{*}$ as $n$ tends to infinity.
$E[V(P)]$ $=\sum_{i}\frac{E[T_{i}]\mu_{i}}{n}$ (12)
$=\sum_{i}\frac{E[T_{i}](\mu^{*}-\Delta_{i})}{n}$ (13)
$=\mu^{*}-\frac{\sum_{i}E[T_{i}]\Delta_{i}}{n}$ (14)
where $\Delta_{i}=\mu^{*}-\mu_{i}$, and $\sum_{i}E[T_{i}]\Delta_{i}$ is the regret $R_{n}$ of the algorithm (see Appendix A).
Since $R_{n}$ growths in a sub-linear rate, so we can get
$\lim_{n\to\infty}E[V(P)]=\mu^{*}-\lim_{n\to\infty}\frac{R_{n}}{n}=\mu^{*}$ (15)
The following theorem provides an upper bound for the growth rate of $R_{n}$ for choosing arm in AlphaZero’s way , thereby demonstrating that the regret is sub-linearly increasing.
Theorem 4.3.
Denote the prior for arm $i$ as $P_{i}$. Without loss of generality, suppose $\mu^{*}=\mu_{1}$. Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of AlphaZero satisfies (Appendix A)
The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear. Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound. Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm. If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to priors. As long as the number of simulations is sufficient, it can still achieve a good target policy.
When reusing information in the search process, we can obtain a reference value for an arm. The existence of this reference value can eliminate the need for exploration of known arms, thereby further reducing the algorithm’s regret. The following theorem proves this point.
Theorem 4.4.
For a bandit which satisfies the assumptions in 4.3 and an already known arm value, the regret upper bound can be reduced if we use the way of 4.2.1 to choose arm. Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (18)
$+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (19)
$+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (20)
The complete proof can be found in Appendix A.
@@ 4.2.2 Theoretical Analysis
MuZero (Schrittwieser et al., 2019) $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
@@
5 Experiment
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. To empirically validate the efficiency of the two boosting techniques introduced in Section 3.2 for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in (Niu et al., 2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces. For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023). The full implementation details available in Appendix B. Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
In the next subsections, we sequentially explore and try to answer the following three questions:
  • What is the concrete effect and hyper-parameter sensitivity of the Just-in-Time Reanalyze technique in ReZero?
  • How much search budgets and time costs can be saved in practice by the Temporal Information Reuse technique?
@@ 5.1 Time Efficiency
Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases). Key parameters are aligned with those from original papers. Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022; Schwarzer et al., 2023) to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix B.
Results Figure 3 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments. It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks. Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 6 and 7. For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman. The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps. The vertical axis indicates Return. Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
@@ 5.2 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the Just-in-Time Reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency. Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 4 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
5.3 Effect of Information Reuse
Indicators Avg. time (ms) tree search (num calls) dynamics model (num calls) data process (num calls)
@@ 5.3 Effect of Information Reuse
Table 2: Comparisons about detailed indicators between ReZero-M and MuZero inside the tree search. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls.
To further validate and understand the advantages and significance of the Information Reuse technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. Comparative analysis between ReZero-M and MuZero reveals that the Information Reuse technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. Besides, Information Reuse also incrementally reallocates computations (tree search simulations) towards previously non-visited nodes rather than repeatedly accessing extant nodes, which maybe yield some performance gains in some environments (e.g. Breakout in Figure 3).
@@ 6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research. These explorations could broadening the application horizons of MCTS-based algorithms.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
8 Acknowledgements
Thanks to each of the authors for their efforts in the work of ReZero. We spent several memorable mid-nights and early mornings together in the lab. Thanks to everyone at OpenDILab and their pioneering work, especially the LightZero benchmark (https://github.com/opendilab/LightZero) for providing the base code for ReZero. Finally, we would like to thank Shanghai AI Laboratory and professor Jing Yang of Xi’an Jiaotong University for providing a platform for this work.

### v3

@@ 4.2 Backward-view Reanalyze
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order. Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$. We denote $S_{l}^{t}$ as the $t$-th state in the $l$-th trajectory. We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on. After we search on state $S_{l}^{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained. When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
$a_{root}=\arg\max_{a}I_{l}^{t}(a)$ (8)
$I_{t}(a)=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (9)
where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model. If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10]. If action $a_{l}^{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version. Notice that the analysis in Proposal 2 assumes that we know the expectation value. However, in our design, we use the root value as an approximate estimation. Still based on the stochastic bandit assumption, the following theorem analyzes the impact of the error between these two values:
Theorem 1
For a bandit which satisfies assumptions in Proposal 1 and a sample result for an arm, the regret upper bound can be reduced if we use the way of Equation 8 to choose arm. Specifically, if the sampled arm is the optimal arm with $n$ times sample mean $\hat{\mu}_{1n}$, then the regret $R_{n}$ satisfies:
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+n\exp(-\frac{n\varepsilon^{2}}{2}))$ (10)
In other case, with the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (11)
$+$ $\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2}))$ (12)
This theorem demonstrates that using the sampled mean as a substitute for the expectation value can also reduce the asymptotic upper bound of regret and maintain sub linear growth rate. This can be understood as follows: although there is an error between the sampled mean and the expectation value, since we use the average of $n$ samples (as the root value comes from a tree search which has $n$ simulations), this sampled mean will increasingly approximate the expectation value as $n$ grows, thereby preventing a linear increase. Refer to Appendix A for the proof.
@@
4.3 The ReZero Framework
The technique introduced in Section 4.2 will bring a new problem in practice. The experiments in the Appendix C.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed. However, as shown in the Figure 2, to conduct the backward-view reanalyze, we need to divide the batch into $\frac{1}{k+1}$ of its original size. This diminishes the benefits of parallelized search and, on the contrary, makes the algorithm slower.
Figure 3: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases. The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution. The model is updated $n$ iterations between two collections. MuZero reanalyzes the mini-batch before each model update. ReZero reanalyzes the entire buffer after certain iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
We propose a new pipeline that is more compatible with the method introduced in Section 4.2. In particular, during the collect phase, we have transitioned from using MCTS to select actions to directly sampling actions based on the policy network’s output. This shift can be interpreted as an alternative method to augment exploration in the collect phase, akin to the noise injection at the root node in MCTS-based algorithms. Figure 10 in Appendix C.3 indicates that this modification does not significantly compromise performance during evaluation. For the reanalyze process, we introduce the periodical entire-buffer reanalyze mechanism. As shown in Figure 3, we reanalyze the whole buffer after a fixed number of training iterations. For each training iteration, we don’t need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN [28], maintaining a constant policy target for a certain number of training iterations. Reducing the frequency of policy target updates correspondingly decreases the number of MCTS calls. Concurrently, this does not result in a decrease in performance. 2 Since we no longer invoke MCTS during the collect phase, all MCTS calls are concentrated in the reanalyze process. And during the entire-buffer reanalyze, we are no longer constrained by the size of the mini-batch, allowing us to freely adjust the batch size to leverage the advantages of large batches. Figure 11 shows both excessively large and excessively small batch sizes can lead to a decrease in search speed. We choose the batch size of $2000$ according to the experiment in Appendix C.3.
Experiments show that our algorithm maintains high sample efficiency and greatly save the running time of the algorithm. Our pipeline also has the following potential improvement directions:
  • When directly using policy for data collection, action selection is no longer bound by tree search. Thus, previous vectorized environments like [15] can be seamlessly integrated. Besides, this design makes MCTS-based algorithms compatible to existing RL exploration methods like [29].
  • We can use a more reasonable way, such as weighted sampling to preferentially reanalyze a part of the samples in the buffer, instead of simply reanalyzing all samples in the entire buffer. This is helpful to further reduce the computational overhead.
5 Experiment
@@
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. To empirically validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in [30]: MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces. For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark [30]. The full implementation details available in Appendix B. Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings. In the next subsections, we sequentially explore and try to answer the following three questions:
  • What is the effect and hyper-parameter sensitivity of the Entire-buffer Reanalyze technique?
  • How much search budgets can be saved in practice by the Backward-view Reanalyze technique?
@@
MuZero [10] $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
@@ 5.1 Time Efficiency
Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases). Key parameters are aligned with those from original papers. Specifically, we set the replay ratio (the ratio between environment steps and training steps) [31, 32] to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) [13] is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix B.
Results Figure 4 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments. It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks. Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 7 and 8. For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
Figure 5: The effect of Reanalyze Frequency in ReZero-M on the MsPacman game. The proper reanalyze frequency can improve time and sample efficiency while obtaining the comparable return with MuZero (reanalyze ratio=1).
@@ 5.2 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the entire-buffer reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency. Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 5 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
5.3 Effect of Backward-view Reanalyze
Indicators Avg.time (ms) tree search (num calls) dynamics (num calls) data process (num calls)
@@ 5.3 Effect of Backward-view Reanalyze
Table 2: Comparisons about detailed time indicators between MuZero and ReZero-M inside the tree search.
To further validate and understand the advantages and significance of the backward-view reanalyze technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls. Comparative analysis between ReZero-M and MuZero reveals that the backward-view reanalyze technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. Additionally, Theorem 1 demonstrates that the backward-view reanalyze can reduce the regret upper bound, which indicates a better search result. Figure 9 in Appendix C.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman. The experimental results reveal that our method not only enhances the speed of individual searches but also improves sample efficiency. This aligns with the theoretical analysis.
@@ 6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X [33], constitutes a fertile avenue for future research. These explorations could broadening the application horizons of MCTS-based algorithms.